# Weak convergence of Bayes estimators under general loss functions

### v1

@@ A.3 Differentiability of the parametric Wasserstein distance
We define $\mathcal{W}_{2}(\mathcal{X})$ as the space of probability measure on $\mathcal{X}$ with finite second moment.
@@
Let $k\in\mathbb{N}$, $\mathcal{X}\subseteq\mathbb{R}^{k}$ be a convex set, let $\Theta\subseteq\mathbb{R}^{d}$ be open and denote the Lebesgue measure on $\mathbb{R}^{k}$ as $\lambda$. Suppose that $\{P_{\vartheta}:\vartheta\in\Theta\}\subset\mathcal{W}_{2}(\mathcal{X})$ is a parametric family of probability measures with $P_{\vartheta}\ll\lambda$ for all $\vartheta\in\Theta$. We denote $p_{\vartheta}$ as the Lebesgue density of $P_{\vartheta}$ and assume that $\vartheta\mapsto p_{\vartheta}(x)$ is differentiable for any $x\in\mathcal{X}$ and $p_{\vartheta}(x)>0$ for any $x\in\mathcal{X}$ and $\vartheta\in\Theta$. For any $t,\vartheta\in\Theta$ denote
$\Phi(t,\vartheta):=\left\{\phi\in L^{1}(P_{t}):W_{2}^{2}(P_{t},P_{\vartheta})=\int_{\mathcal{X}}\phi(x)\mathop{}\!\mathrm{d}P_{t}(x)+\int_{\mathcal{X}}\phi^{c}(y)\mathop{}\!\mathrm{d}P_{\vartheta}(y)\right\},\quad t,\vartheta\in\Theta,$
as the set of solutions of the dual formulation, where the c-transform is
$\phi^{c}(y)=\inf_{x\in\mathcal{X}}\left\lVert x-y\right\rVert^{2}-\phi(x).$
For $t\in\Theta$ and $\delta>0$ let
@@
Furthermore, for some fixed $t_{0}\in\Theta$ and for any $t,\vartheta\in\Theta$ we impose the following regularity conditions:
@@
    For any $\phi_{t}^{\vartheta}\in\Phi(t,\vartheta)$ there exists $\phi_{t_{0}}^{\vartheta}\in\Phi(t_{0},\vartheta)$ such that
    $\phi_{t}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\longrightarrow 0,$
    as $t\to t_{0}$, for $\lambda$-a.e. $x\in\mathcal{X}$.
@@
    For any $\phi_{t_{0}}^{\vartheta}\in\Phi(t_{0},\vartheta)$ there exists $\delta>0$ such that
    $\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\mathcal{X}$}}$}}\nolimits\sup_{t_{1},t_{2}\in B_{\delta}(t_{0})}\left\lvert\phi_{t_{1}}^{\vartheta}(x)\frac{\partial}{\partial t_{i}}p_{t}(x)\bigg|_{t=t_{2}}\right\rvert\mathop{}\!\mathrm{d}x<\infty$
    for all $i=1,\dots,d$.
@@
$\lim\limits_{h\to 0}I_{i}(h)=\frac{\partial}{\partial t_{i}}W_{2}^{2}(P_{t},P_{\vartheta})|_{t=t_{0}}.$
Recall the dual formulation of the squared 2-Wasserstein distance (Villani,, 2009, Theorem 5.10). For any $\mu,\nu\in\mathcal{W}_{2}(\mathcal{X})$ it holds
@@
where for any $\phi\in L^{1}(\mu)$
$\phi^{c}(y)=\inf_{x\in\mathcal{X}}\left\lVert x-y\right\rVert^{2}-\phi(x).$
For $t,\vartheta\in\Theta$ denote $\phi_{t}^{\vartheta}\in\Phi(t,\vartheta)$ as any solution to the dual problem. By Villani, (2009, Theorem 5.10) the set $\Phi(t,\vartheta)$ is nonempty. Then, we have for any $i=1,\dots,d$, $\phi_{t_{0}+e_{i}h}^{\vartheta}\in\Phi(t_{0}+e_{i}h,\vartheta)$ and $\phi_{t_{0}}^{\vartheta}\in\Phi(t_{0},\vartheta)$
@@
By Assumption (i) for any $i=1,\dots,d$ there exists $\phi_{t_{0}}^{\vartheta}(x)\in\Phi(t_{0},\vartheta)$ such that
$\phi_{t_{0}+e_{i}h}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\to 0,\quad h\to 0,$
for $\lambda$-a.e. $x\in\mathcal{X}$. The mean value theorem implies that for fixed $x\in\mathcal{X}$, $t_{0}\in\Theta$ and for $\left\lvert h\right\rvert\leq\delta$ it holds
$\left\lvert\frac{p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)}{h}\right\rvert=\left\lvert\frac{\partial}{\partial t_{i}}p_{t}(x)\bigg|_{t=\tilde{t}}\right\rvert$
for some $\tilde{t}=t_{0}+e_{i}h_{0}$ and $\left\lvert h_{0}\right\rvert\leq\delta$. Denote $B_{\delta}(t_{0})=\{s\in\Theta:\left\lVert s-t_{0}\right\rVert\leq\delta\}$. For any fixed $x\in\mathcal{X}$ we obtain the upper bound
$\left\lvert\left(\phi_{t+e_{i}h}^{\vartheta}(x)-\phi_{t}^{\vartheta}(x)\right)\frac{p_{t+e_{i}h}(x)-p_{t}(x)}{h}\right\rvert$ $\leq\sup_{t_{1},t_{2}\in B_{\delta}(t_{0})}\left\lvert\left(\phi_{t_{1}}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\right)\frac{\partial}{\partial t_{i}}p_{t}(x)\bigg|_{t=t_{2}}\right\rvert$
$\leq 2\sup_{t_{1},t_{2}\in B_{\delta}(t_{0})}\left\lvert\phi_{t_{1}}^{\vartheta}(x)\frac{\partial}{\partial t_{i}}p_{t}(x)\bigg|_{t=t_{2}}\right\rvert$.
By Assumption (ii) we can choose $\delta>0$ small enough such that the upper bound is integrable. Analogously, by Assumption (ii) we obtain the integrable upper bound
$\left\lvert\phi_{t_{0}}^{\vartheta}(x)\frac{p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)}{h}\right\rvert\leq\sup_{t_{1},t_{2}\in B_{\delta}(t_{0})}\left\lvert\phi_{t_{1}}^{\vartheta}(x)\frac{\partial}{\partial t_{i}}p_{t}(x)\bigg|_{t=t_{2}}\right\rvert$.
Consequently, Lebesgue’s dominated convergence theorem and differentiability of the map $\vartheta\mapsto p_{\vartheta}(x)$ imply
@@
$\leq\lim\limits_{h\to 0}\left(\int_{\mathcal{X}}\frac{\phi_{t_{0}}^{\vartheta}(x)\left(p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)\right)}{h}\mathop{}\!\mathrm{d}x+\int_{\mathcal{X}}\frac{\left(\phi_{t_{0}+e_{i}h}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\right)\left(p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)\right)}{h}\mathop{}\!\mathrm{d}x\right)$
$=\int_{\mathcal{X}}\phi_{t_{0}}^{\vartheta}(x)\frac{\partial}{\partial t_{i}}p_{t}(x)|_{t=t_{0}}\mathop{}\!\mathrm{d}x.$
@@
$\liminf_{h\to 0}I_{i}(h)\geq\int_{\mathcal{X}}\phi_{t_{0}}^{\vartheta}(x)\frac{\partial}{\partial t_{i}}p_{t}(x)|_{t=t_{0}}\mathop{}\!\mathrm{d}x.$
@@
Assumption (i) is a continuity condition on the dual solutions in the parameter. By Staudt et al., (2025, Theorem 2), for the setting of the previous Theorem, the dual solutions are unique up to additive constants. Moreover, Assumption (ii) implies that
$\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\mathcal{X}$}}$}}\nolimits\sup_{t\in B_{\delta}(t_{0})}\left\lvert\frac{\partial}{\partial t_{i}}p_{t}(x)\right\rvert\mathop{}\!\mathrm{d}x<\infty$
@@
Hence, the formula of the gradient of the Wasserstein distance is independent of the choice of the potential $\phi_{t_{0}}^{\vartheta}\in\Phi(t_{0},\vartheta)$.
If the space $\mathcal{X}$ is a subset of the real line, we obtain a more concrete form for the second derivative and sufficient conditions for differentiability of the map $t\mapsto\nabla_{t}W_{2}^{2}(P_{t},P_{\vartheta})$. The next result is inspired by Theorem 1 from Chen and Li, (2020), but does not assume that $\mathcal{X}$ is compact. We denote $\overline{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$ as the extended real numbers.
@@
Let $T_{t}^{\vartheta}(x)$ denote the optimal transport map transporting $P_{t}$ to $P_{\vartheta}$. For $t\in\Theta$ and $\delta>0$ define
@@
    $\int_{\mathcal{X}}\sup_{t\in B_{\delta}(t_{0})}\left\lvert\frac{\partial}{\partial t_{i}}p_{t}(x)\right\rvert\mathop{}\!\mathrm{d}x<\infty$
    for all $i=1,\dots,d$.
@@
    The map $\vartheta\mapsto F_{\vartheta}(x)$ is twice differentiable for any $x\in\mathcal{X},\vartheta\in\Theta$ and for any $\vartheta\in\Theta$ there exists $\eta>0$ such that
    $\int_{\mathcal{X}}\sup_{t\in B_{\eta}(t_{0})}\left\lvert(T_{t}^{\vartheta}(x)-x)\left(\frac{\partial^{2}}{\partial t_{i}t_{j}}F_{t}(x)\right)+\left(\frac{\partial}{\partial t_{i}}T_{t}^{\vartheta}(x)\right)\left(\frac{\partial}{\partial t_{j}}F_{t}(x)\right)\right\rvert\mathop{}\!\mathrm{d}x<\infty$
    for all $i,j=1,\dots,d$.
@@
    For any $\vartheta\in\Theta$ and $x\in\mathcal{X}$ it holds
    $\lim\limits_{x\to b}\nabla_{t}F_{t}(x)|_{t=t_{0}}\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911pta$}}^{\text{\normalsize$\scriptstyle x$}}$}}\nolimits y-T_{t_{0}}^{\vartheta}(y)\mathop{}\!\mathrm{d}y=0.$
If the map $t\mapsto W_{2}^{2}(P_{t},P_{\vartheta})$ is differentiable for any $\vartheta\in\Theta$, then the second derivative at $t_{0}$ exists and it holds
@@
From Theorem 4, we know that the first derivative is given by
@@
By Gangbo and McCann, (1996, Theorem 1.2) it holds
for $P_{t_{0}}$-a.e. $x\in\mathcal{X}$. As the density of $P_{t_{0}}$ is strictly positive on $\mathcal{X}$, we conclude that this also holds $\lambda$-a.e. $x\in\mathcal{X}$. Hence, for some constant $C\in\mathbb{R}$
$\phi_{t_{0}}^{\vartheta}(x)=\int_{a}^{x}2(y-T_{t_{0}}^{\vartheta}(y))\mathop{}\!\mathrm{d}y+C.$ (8)
Furthermore, by Villani, (2003, Theorem 2.18) the optimal transport map is given by $T_{t_{0}}^{\vartheta}(x)=F^{-1}_{\vartheta}(F_{t_{0}}(x))$. Integration by parts yields
@@
By Assumption (i) we are allowed to interchange order of derivation and integration, which leads to
$\lim\limits_{x\to b}C\nabla_{t}F_{t}(x)|_{t=t_{0}}=\lim\limits_{x\to b}C\int_{a}^{x}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=C\int_{\mathcal{X}}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=\nabla_{t}\left(\int_{\mathcal{X}}p_{t}(y)\mathop{}\!\mathrm{d}y\right)|_{t=t_{0}}=0.$
Here we used that $p_{t}$ integrates to 1. If we combine this fact and equation (8) with Assumption (iii), it follows
$\lim\limits_{x\to b}\phi_{t_{0}}^{\vartheta}(x)\nabla_{t}F_{t}(x)|_{t=t_{0}}$ $=\lim\limits_{x\to b}C\nabla_{t}F_{t}(x)|_{t=t_{0}}+\nabla_{t}F_{t}(x)|_{t=t_{0}}\int_{a}^{x}2\left(y-F_{\vartheta}^{-1}\left(F_{t_{0}}(y)\right)\right)\mathop{}\!\mathrm{d}y$
$=0.$
Furthermore, it holds
$\lim\limits_{x\to a}\phi_{t_{0}}^{\vartheta}(x)=\lim\limits_{x\to a}\int_{a}^{x}2(y-T_{t_{0}}^{\vartheta}(y))dy+C=C.$
Again, by the dominated convergence theorem we obtain
$\lim\limits_{x\to a}\int_{a}^{x}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=0$
@@
$\mathbb{I}(a<y\leq x)\nabla_{t}p_{t}(y)|_{t=t_{0}}\leq\nabla_{t}p_{t}(y)|_{t=t_{0}}$
and consequently
$\lim\limits_{x\to a}\phi_{t_{0}}^{\vartheta}(x)\nabla_{t}F_{t}(x)|_{t=t_{0}}=\lim\limits_{x\to a}\left(\int_{a}^{x}2(y-T_{t_{0}}^{\vartheta}(y))\mathop{}\!\mathrm{d}y+C\right)\int_{a}^{x}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=0.$
That means we have derived that
@@
The derivative with respect to $x$ becomes
@@
On the other hand taking the derivative with respect to $t$ at $t=t_{0}$ yields
@@
Recall that Assumption (ii) enables the dominated convergence theorem, c.f. Theorem 4, so that we are allowed to interchange integral and derivative. This leads to

### v2

@@ A.3 Differentiability of the parametric Wasserstein distance
Let $\mathcal{W}_{2}(\mathcal{X})$ be the space of probability measure on $\mathcal{X}$ with finite second moment. For a parametric family of probability measures $\{P_{\vartheta}:\vartheta\in\Theta\}\subset\mathcal{W}_{2}(\mathcal{X})$ that are absolutely continuous with respect to the Lebesgue measure, we denote, for any $t,\vartheta\in\Theta$,
$\Phi(t,\vartheta):=\left\{\phi\in L^{1}(P_{t}):W_{2}^{2}(P_{t},P_{\vartheta})=\int_{\mathcal{X}}\phi(x)\mathop{}\!\mathrm{d}P_{t}(x)+\int_{\mathcal{X}}\phi^{c}(y)\mathop{}\!\mathrm{d}P_{\vartheta}(y)\right\}$
as the set of solutions to the dual formulation (see e.g. Santambrogio,, 2015), where the $c$-transform is
$\phi^{c}(y)=\inf_{x\in\mathcal{X}}\left\lVert x-y\right\rVert^{2}-\phi(x).$ (8)
The set $\Phi(t,\vartheta)$ is nonempty by (Villani,, 2009, Theorem 5.10).
@@
Let $k\in\mathbb{N}$, $\mathcal{X}\subseteq\mathbb{R}^{k}$ be a convex set, let $\Theta\subseteq\mathbb{R}^{d}$ be open and denote the Lebesgue measure on $\mathbb{R}^{k}$ as $\lambda$. Suppose that $\{P_{\vartheta}:\vartheta\in\Theta\}\subset\mathcal{W}_{2}(\mathcal{X})$ is a parametric family of probability measures with $P_{\vartheta}\ll\lambda$ for all $\vartheta\in\Theta$. We denote $p_{\vartheta}$ as the Lebesgue density of $P_{\vartheta}$ and assume that $\vartheta\mapsto p_{\vartheta}(x)$ is differentiable for any $x\in\mathcal{X}$ and $p_{\vartheta}(x)>0$ for any $x\in\mathcal{X}$ and $\vartheta\in\Theta$. For $t\in\Theta$ and $\delta>0$, let
@@
Furthermore, for some fixed $t_{0}\in\Theta$ and for any $t,\vartheta\in\Theta$, we assume that there exists a family of nonempty subsets $\Phi_{0}(t,\vartheta)\subseteq\Phi(t,\vartheta)$ such that:
@@
    For any sequence $t_{n}\to t_{0}$, there exists a sequence $\phi_{t_{n}}^{\vartheta}\in\Phi_{0}(t_{n},\vartheta)$ and a fixed element $\phi_{t_{0}}^{\vartheta}\in\Phi_{0}(t_{0},\vartheta)$ such that, as $n\to\infty$, for $\lambda$-a.e. $x\in\mathcal{X}$,
    $\phi_{t_{n}}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\longrightarrow 0.$
@@
    There exists $\delta>0$ such that, for all $i=1,\dots,d$,
    $\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\mathcal{X}$}}$}}\nolimits\sup_{t,\tilde{t}\in B_{\delta}(t_{0})}\sup_{\phi\in\Phi_{0}(t,\vartheta)}\left\lvert\phi(x)\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert\mathop{}\!\mathrm{d}x<\infty.$
@@
$\lim\limits_{h\to 0}I_{i}(h)={\frac{\partial}{\partial v_{i}}W_{2}^{2}(P_{v},P_{\vartheta})\Big|_{v=t_{0}}}.$
Recall the dual formulation of the squared 2-Wasserstein distance (Villani,, 2009, Theorem 5.10). For any $\mu,\nu\in\mathcal{W}_{2}(\mathcal{X})$, it holds
@@
where $\phi^{c}$ is the $c$-transform of $\phi$, defined in (8). By condition (i), for any $i=1,\ldots,d$, there exist a sequence $\phi_{t_{0}+e_{i}h}^{\vartheta}\in\Phi_{0}(t_{0}+e_{i}h,\vartheta)\subset\Phi(t_{0}+e_{i}h,\vartheta)$ and a fixed $\phi_{t_{0}}^{\vartheta}\in\Phi_{0}(t_{0},\vartheta)$ such that
$\phi_{t_{0}+e_{i}h}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\to 0,\quad h\to 0,$ (9)
for $\lambda$-a.e. $x\in\mathcal{X}$. By the dual formulation, we obtain
$W_{2}^{2}(P_{t_{0}},P_{\vartheta})\geq\int_{\mathcal{X}}\frac{\phi_{t_{0}+e_{i}h}^{\vartheta}(x)p_{t_{0}}(x)}{h}\mathop{}\!\mathrm{d}x+\int_{\mathcal{X}}\frac{\left(\phi_{t_{0}+e_{i}h}^{\vartheta}\right)^{c}(x)p_{\vartheta}(x)}{h}\mathop{}\!\mathrm{d}x$
and consequently
@@
The mean value theorem implies that for fixed $x\in\mathcal{X}$, $t_{0}\in\Theta$ and for $\left\lvert h\right\rvert\leq\delta$ it holds
$\left\lvert\frac{p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)}{h}\right\rvert={\left\lvert\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert}$
for some $\tilde{t}=t_{0}+e_{i}h_{0}$ and $\left\lvert h_{0}\right\rvert\leq\delta$. Denote $B_{\delta}(t_{0})=\{s\in\Theta:\left\lVert s-t_{0}\right\rVert\leq\delta\}$. For any fixed $x\in\mathcal{X}$, we obtain the upper bound
$\left\lvert\left(\phi_{t_{0}+e_{i}h}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\right)\frac{p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)}{h}\right\rvert\leq\sup_{t,\tilde{t}\in B_{\delta}(t_{0})}{\sup_{\phi\in\Phi_{0}(t,\vartheta)}\left\lvert\left(\phi(x)-\phi_{t_{0}}^{\vartheta}(x)\right)\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert}$
$\leq{\sup_{t,\tilde{t}\in B_{\delta}(t_{0})}\sup_{\phi\in\Phi_{0}(t,\vartheta)}\left\lvert\phi(x)\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert+\sup_{\tilde{t}\in B_{\delta}(t_{0})}\left\lvert\phi^{\vartheta}_{t_{0}}(x)\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert}$
$\leq 2\sup_{t,\tilde{t}\in B_{\delta}(t_{0})}{\sup_{\phi\in\Phi_{0}(t,\vartheta)}\left\lvert\phi(x)\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert}$.
By condition (ii), we can choose $\delta>0$ small enough such that the upper bound is integrable. Analogously, we obtain the following upper bound that is integrable,
$\left\lvert\phi_{t_{0}}^{\vartheta}(x)\frac{p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)}{h}\right\rvert\leq\sup_{t,\tilde{t}\in B_{\delta}(t_{0})}{\sup_{\phi\in\Phi_{0}(t,\vartheta)}\left\lvert\phi(x)\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=\tilde{t}}\right\rvert}$.
Consequently, Lebesgue’s dominated convergence theorem and differentiability of the mapping $\vartheta\mapsto p_{\vartheta}(x)$ imply
@@
$\leq$ $\lim\limits_{h\to 0}\Biggl[\int_{\mathcal{X}}\frac{\phi_{t_{0}}^{\vartheta}(x)\left(p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)\right)}{h}\mathop{}\!\mathrm{d}x+\int_{\mathcal{X}}\frac{\left(\phi_{t_{0}+e_{i}h}^{\vartheta}(x)-\phi_{t_{0}}^{\vartheta}(x)\right)\left(p_{t_{0}+e_{i}h}(x)-p_{t_{0}}(x)\right)}{h}\mathop{}\!\mathrm{d}x\Biggr]$
$=$ $\int_{\mathcal{X}}\phi_{t_{0}}^{\vartheta}(x){\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=t_{0}}}\mathop{}\!\mathrm{d}x.$
@@
$\liminf_{h\to 0}I_{i}(h)\geq\int_{\mathcal{X}}\phi_{t_{0}}^{\vartheta}(x){\frac{\partial}{\partial v_{i}}p_{v}(x)\Big|_{v=t_{0}}}\mathop{}\!\mathrm{d}x.$
@@
In Theorem 4, condition (i) is a continuity condition on the dual solutions in the parameter. By Staudt et al., (2025, Theorem 2), the dual solutions are unique up to additive constants. Moreover, condition (ii) implies that
$\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\mathcal{X}$}}$}}\nolimits{\sup_{v\in B_{\delta}(t_{0})}\left\lvert\frac{\partial}{\partial v_{i}}p_{v}(x)\right\rvert}\mathop{}\!\mathrm{d}x<\infty$
@@
Hence, the formula of the gradient of the Wasserstein distance is independent of the choice of the concrete potential and it holds
$\nabla_{t}W_{2}^{2}(P_{t},P_{\vartheta})|_{t=t_{0}}=\int_{\mathcal{X}}\phi(x)\nabla_{t}p_{t}(x)|_{t=t_{0}}\mathop{}\!\mathrm{d}x=\int_{\mathcal{X}}\psi(x)\nabla_{t}p_{t}(x)|_{t=t_{0}}\mathop{}\!\mathrm{d}x$
for any $\phi,\psi\in\Phi(t_{0},\vartheta)$. The derivative of the squared 2-Wasserstein distance depends on $\vartheta$ only through $\phi_{t_{0}}^{\vartheta}\in\Phi(t_{0},\vartheta)$. The collection of subsets $\{\Phi_{0}(t,\vartheta):t\in\Theta\}$ is typically chosen so that each set consists of potentials in $\Phi(t,\vartheta)$ with a fixed additive constant. In this case, $\Phi_{0}(t,\vartheta)$ is a singleton for every $t,\vartheta\in\Theta$.
If the space $\mathcal{X}$ is a subset of the real line, we obtain a more concrete form for the second derivative and sufficient conditions for differentiability of the mapping $t\mapsto\nabla_{t}W_{2}^{2}(P_{t},P_{\vartheta})$. The next result is inspired by Theorem 1 from Chen and Li, (2020), but does not assume that $\mathcal{X}$ is compact. We denote $\overline{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$ as the extended real numbers.
@@
Let $T_{t}^{\vartheta}(x)$ denote the optimal transport map from $P_{t}$ to $P_{\vartheta}$. For $t\in\Theta$ and $\delta>0$ define
@@
    $\int_{\mathcal{X}}{\sup_{v\in B_{\delta}(t_{0})}\left\lvert\frac{\partial}{\partial v_{i}}p_{v}(x)\right\rvert}\mathop{}\!\mathrm{d}x<\infty$
    for all $i\in\{1,\dots,d\}$.
@@
    The map $\vartheta\mapsto F_{\vartheta}(x)$ is twice differentiable for any $x\in\mathcal{X},\vartheta\in\Theta$, and for any $\vartheta\in\Theta$ there exists $\eta>0$ such that
    $\int_{\mathcal{X}}{\sup_{v\in B_{\eta}(t_{0})}\left\lvert\left(T_{v}^{\vartheta}(x)-x\right)\left(\frac{\partial^{2}}{\partial v_{i}v_{j}}F_{v}(x)\right)+\left(\frac{\partial}{\partial v_{i}}T_{v}^{\vartheta}(x)\right)\left(\frac{\partial}{\partial v_{j}}F_{v}(x)\right)\right\rvert}\mathop{}\!\mathrm{d}x<\infty$
    for all $i,j\in\{1,\dots,d\}$.
@@
    For any $\vartheta\in\Theta$ and $x_{0}\in\mathcal{X}$, it holds
    $\lim\limits_{x\to a}\nabla_{t}F_{t}(x)|_{t=t_{0}}\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911ptx_{0}$}}^{\text{\normalsize$\scriptstyle x$}}$}}\nolimits\left(y-T_{t_{0}}^{\vartheta}(y)\right)\mathop{}\!\mathrm{d}y=0,\;\lim\limits_{x\to b}\nabla_{t}F_{t}(x)|_{t=t_{0}}\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911ptx_{0}$}}^{\text{\normalsize$\scriptstyle x$}}$}}\nolimits\left(y-T_{t_{0}}^{\vartheta}(y)\right)\mathop{}\!\mathrm{d}y=0.$
If the mapping $t\mapsto W_{2}^{2}(P_{t},P_{\vartheta})$ is differentiable for any $\vartheta\in\Theta$, with the derivative
for any $\phi_{t_{0}}^{\vartheta}\in\Phi(t_{0},\vartheta)$, then the second derivative at $t_{0}$ exists, and it holds
@@
By Villani, (2003, Theorem 2.18), the optimal transport map is given by $T_{t_{0}}^{\vartheta}(x)=F^{-1}_{\vartheta}(F_{t_{0}}(x))$. By Gangbo and McCann, (1996, Theorem 1.2), it holds, for any $\phi_{t_{0}}^{\vartheta}(x)\in\Phi(t_{0},\vartheta)$,
for $P_{t_{0}}$-a.e. $x\in\mathcal{X}$. As the density of $P_{t_{0}}$ is strictly positive on $\mathcal{X}$, we conclude that (10) also holds $\lambda$-a.e. $x\in\mathcal{X}$. Hence, for any $C\in\mathbb{R}$ and $x_{0}\in\mathcal{X}$, the function
$x\mapsto\int_{x_{0}}^{x}2\left(y-T_{t_{0}}^{\vartheta}(y)\right)\mathop{}\!\mathrm{d}y+C,$} (11)
is included in the set $\Phi(t_{0},\vartheta)$. Fix some $C\in\mathbb{R}$ and $x_{0}\in\mathcal{X}$ and set
$\phi_{t_{0}}^{\vartheta}(x)=\int_{x_{0}}^{x}2\left(y-T_{t_{0}}^{\vartheta}(y)\right)\mathop{}\!\mathrm{d}y+C\;\in\;\Phi(t_{0},\vartheta).$
The first derivative is given by
@@
Integration by parts yields
@@
By (11) and condition (iii), it holds
$\lim\limits_{x\to b}\phi_{t_{0}}^{\vartheta}(x)\nabla_{t}F_{t}(x)|_{t=t_{0}}$ $=\lim\limits_{x\to b}\left(C\nabla_{t}F_{t}(x)|_{t=t_{0}}+\nabla_{t}F_{t}(x)|_{t=t_{0}}\int_{x_{0}}^{x}2\left(y-F_{\vartheta}^{-1}\left(F_{t_{0}}(y)\right)\right)\mathop{}\!\mathrm{d}y\right)$
$=\lim\limits_{x\to b}C\nabla_{t}F_{t}(x)|_{t=t_{0}}.$
By condition (i), we can interchange order of derivation and integration, which leads to
$\lim\limits_{x\to b}C\nabla_{t}F_{t}(x)|_{t=t_{0}}=\lim\limits_{x\to b}C\int_{a}^{x}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=C\int_{\mathcal{X}}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=\nabla_{t}\left(\int_{\mathcal{X}}p_{t}(y)\mathop{}\!\mathrm{d}y\right)\Big|_{t=t_{0}}=0.$
Here we used that $p_{t}$ integrates to 1. This shows that $\lim\limits_{x\to b}\phi_{t_{0}}^{\vartheta}(x)\nabla_{t}F_{t}(x)|_{t=t_{0}}=0$. By the dominated convergence theorem, we obtain
$\lim\limits_{x\to a}\nabla_{t}F_{t}(x)|_{t=t_{0}}=\lim\limits_{x\to a}\int_{a}^{x}\nabla_{t}p_{t}(y)|_{t=t_{0}}\mathop{}\!\mathrm{d}y=0$
@@
$\mathbb{I}(a<y\leq x)\nabla_{t}p_{t}(y)|_{t=t_{0}}\leq\nabla_{t}p_{t}(y)|_{t=t_{0}}.$
By this convergence and condition (iii), it holds
$\lim\limits_{x\to a}\phi_{t_{0}}^{\vartheta}(x)\nabla_{t}F_{t}(x)|_{t=t_{0}}=\lim\limits_{x\to a}\nabla_{t}F_{t}(x)|_{t=t_{0}}\int_{x_{0}}^{x}2\left(y-T_{t_{0}}^{\vartheta}(y)\right)\mathop{}\!\mathrm{d}y+C\nabla_{t}F_{t}(x)|_{t=t_{0}}=0.$
Thus, we have derived that
@@
The derivative with respect to $x$ is
@@
On the other hand, taking the derivative with respect to $t$ at $t=t_{0}$ yields
@@
Recall that condition (ii) enables the dominated convergence theorem, c.f. Theorem 4, so we are allowed to interchange integral and derivative. This leads to